# Concept-TRAK: Understanding how diffusion models learn concepts through concept-level attribution

### v1

@@ 1 Introduction
Data attribution methods [22, 29, 14], which estimate how much each training example contributes to a generated output, have become valuable tools for tasks such as data valuation [20], curation [26] and understanding model behavior [35]. The impressive success of large-scale diffusion models [38, 42, 17] in image generation, often coupled with substantial commercial impact [31, 10, 34], has been largely driven by access to massive training datasets [36]. However, this dependence raises pressing societal concerns, especially around copyright and data ownership [50, 40, 5, 39, 13, 4]. A key question is whether generated outputs can be traced back to specific training samples, enabling the detection of copyrighted content used during training—an area where data attribution methods show significant promise [9].
While recent work has begun exploring attribution methods tailored for diffusion models [52, 23, 27, 48], these approaches generally estimate contributions at the level of entire images. This broad perspective poses a critical limitation: in many practical scenarios, stakeholders care about specific concepts within an image, rather than the whole composition.
@@
To address this, we introduce the notion of concept-level attribution, which estimates each training example’s contribution to specific semantic features such as styles, objects, or concepts. Building on this, we propose Concept-TRAK, an extension of influence functions [22], to quantify how training data affects the model’s output probabilities for individual concepts. Concept-TRAK incorporates two key innovations: (1) a reformulated diffusion training loss based on diffusion posterior sampling (DPS), yielding robust, sample-specific attribution gradients; and (2) a concept-aware reward function that emphasizes semantic relevance. These components significantly enhance attribution fidelity. As shown in Figure 1(b), Concept-TRAK correctly identifies training samples responsible for the concept of Pikachu, rather than irrelevant stylistic cues—enabling precise, interpretable, and stakeholder-relevant attribution for tasks such as copyright auditing and model transparency.
To rigorously evaluate our method, we adopt a retrieval-based benchmark, AbC benchmark [47], for concept attribution. In this setting, Concept-TRAK significantly outperforms prior methods by accurately retrieving training examples that influence specific concepts. Moreover, we demonstrate its practical value through case studies that: (1) identify training samples that contributed most to learning IP-protected concepts, (2) detect the sources of learning unsafe concepts, (3) trace the origins of both desirable features and problematic outputs for model debugging, and (4) provide insights into the origins of relational knowledge (e.g., shake hands, hugs) in diffusion models. Together, these results highlight how concept-level attribution offers valuable insights both for practical applications and for advancing our understanding of how generative models learn and combine concepts from training data.
@@ 2.1 Diffusion Model
Diffusion models [38, 42, 17] are a class of generative models that synthesize images through an iterative denoising process. Starting from a clean image $x_{0}$, the forward process adds Gaussian noise to produce a sequence of increasingly noisy images $x_{t}$, following, $q(x_{t}\mid x_{0})=\mathcal{N}(\sqrt{\alpha_{t}}x_{0},(1-\alpha_{t})I)$, where $\alpha_{t}$ is a noise schedule controlling the level of corruption at timestep $t$.
@@
Diffusion models can be extended to conditional generation by incorporating additional information $c$, such as a text prompt [16, 28]. In this setting, the model learns $\epsilon_{\theta}(x_{t},t;c)\propto\nabla\log p_{t}(x_{t}\mid c)$, allowing it to generate images that are not only realistic but also aligned with the conditioning input.
@@ 2.2 Data Attribution
The goal of data attribution is to estimate the contribution of a training sample $x^{i}$ to a model’s utility loss $\mathcal{V}$—a performance metric or objective function that quantifies how well the model performs (e.g., test loss) [22, 29, 14]. A straightforward approach is the Leave-One-Out method, which involves removing a training example $x^{i}$, retraining the model, and measuring the change in performance. However, this naive strategy is computationally expensive for modern models and large-scale datasets, as it requires retraining the model once for every training point, making it practically infeasible.
To address this limitation, influence functions [22] efficiently approximate the effect of removing a training example $x^{i}$ by using gradient-based estimates, avoiding the need for retraining. Given a model with parameters $\theta\in\mathbb{R}^{d}$ and training loss $\mathcal{L}(\cdot;\theta)$, the influence function is defined as:
@@
Due to the high cost of computing the full Hessian, influence function methods typically rely on inverse Hessian-vector products or approximations such as low-rank factorizations (e.g., EK-FAC) [15]. Despite these advances, influence functions remain challenging to apply in practice. Identifying training samples that strongly impact a target utility [6] requires computing gradients for every sample—an effort equivalent to a full training epoch. Unless these gradients are precomputed and stored, each attribution query results in prohibitive computational cost—especially for today’s large-scale pretrained models trained on massive datasets at great expense. Moreover, storing the full gradient for every training sample is infeasible, since each gradient matches the dimensionality of the model parameters.
To address this, TRAK [29] proposes projecting the gradients into a lower-dimensional space using a random projection matrix $P\in\mathbb{R}^{d\times k}$ with $k\ll d$, and then computing influence in the projected space:
@@ 2.3 Data attribution for diffusion models
Prior work on diffusion models [51, 52, 23] has primarily focused on whole-image attribution, typically using the same objective for both training and utility losses. These studies reveal that attribution performance is highly sensitive to the choice of loss function. In particular, the standard DSM loss introduces stochasticity via both the noise term $\epsilon$ and the perturbed input $x_{t}$, resulting in noisy gradient estimates that require extensive averaging to be reliable—making it suboptimal for attribution. To mitigate this, D-TRAK [52] employs the squared $\ell_{2}$-norm $\|\epsilon_{\theta}\|_{2}^{2}$ to compute influence scores, achieving improved stability and accuracy. However, D-TRAK does not correspond to a proper learning objective tailored to a specific sample $x^{i}_{0}$. Consequently, the resulting gradients may fail to faithfully capture the unique contribution of $x^{i}_{0}$ to the model’s learned behavior.
@@ 3.1 Train Gradients: DPS-Loss
To address the challenges associated with the stochasticity of the DSM loss and the absence of sample-specific attribution for $x^{i}_{0}$, as discussed in Section 2.3, we introduce a novel training objective grounded in the principles of diffusion posterior sampling (DPS) [7], denoted by $\mathcal{L}_{\text{DPS}}$. This objective produces sample-specific training gradients without relying on a fully stochastic regression target and is explicitly designed to capture the influence of upweighting an individual sample $x^{i}_{0}$ on the model’s output.
@@
Consider the effect of finetuning a pretrained diffusion model $\epsilon_{\theta}$ on a single sample $x^{i}_{0}$. To model this process, we adopt a generalized Bayesian update with power likelihood where the influence of $x^{i}_{0}$ is amplified through multiplicative weighting [18] (Appendix B):
@@
Instead of directly training on $x^{i}_{0}$, could we learn to model $\tilde{p}(x)$? One approach would be to use explicit score matching (ESM) [43]. If we had access to a diffusion model $\tilde{\epsilon}$ that already trained on $\tilde{p}(x)$, we could replace DSM with ESM for model training [19]:
@@
A key advantage of ESM is its reduced reliance on the Monte Carlo-estimated regression target used in the DSM loss for influence function computation, resulting in more robust gradient estimates. Moreover, as we will show, $\tilde{\epsilon}$ can incorporate $x^{i}_{0}$ as guidance information—unlike prior approaches [52]. However, directly obtaining $\tilde{\epsilon}$ would require expensive retraining for each sample $x^{i}_{0}$, which is impractical. To address this, we introduce an efficient, feasible, and robust approximation to $\tilde{\epsilon}$.
@@
The core idea of our method is to approximate $\tilde{\epsilon}$ without additional training by using diffusion posterior sampling (DPS) principles [7]. For our target distribution $\tilde{p}(x)\propto p(x)\cdot p(x^{i}_{0}\mid x)^{w}$ , the score function at timestep $t$ becomes:
@@
where $\text{sg}[\cdot]$ denotes stop-gradient. For clarity, we use a slight abuse of notation and denote $\mathcal{L}(x_{t},\theta)$ as the loss at a single timestep $t$.
@@ 3.2 Utility Gradient: Reward-DPS-Loss
Specifically, we leverage the insight that training a model with a reward function $R(x)$ modifies the learned distribution to $\tilde{p}(x)=\frac{1}{Z}p(x)\exp(\frac{1}{\beta}R(x))$, where $Z$ is a normalizing constant and $\beta$ controls the KL-divergence between $p$ and $\tilde{p}$ [33]. For this distribution, the score function at timestep $t$ becomes:
@@
where $\text{sg}[\cdot]$ denotes stop-gradient.
@@
We focus on measuring utility gradients for textual concepts. Extensions to other reward types, including implicit rewards [33], explicit reward models, and rewards for local concept attribution, are discussed in Appendix C.
@@
For textual concepts, we model the reward using conditional probabilities: $R(x_{t})\propto\log p_{t}(c\mid x_{t})$. The gradient of this reward becomes $\nabla_{x_{t}}R(x_{t})\propto\nabla_{x_{t}}\log p_{t}(x_{t}\mid c)-\nabla_{x_{t}}\log p_{t}(x_{t})$, which directly corresponds to the classifier-free guidance [16].
To effectively isolate specific textual concepts, recent research [12, 49, 3, 24] demonstrates the importance of contrastive prompt engineering. Rather than using a single concept descriptor, defining concepts through comparison between positive and negative expressions yields more precise concept isolation. For example, attributing the concept “old” is more effective by comparing “old woman” to “young woman,” as this isolates age within a consistent context. Thus, we define our concept trio: $c$ = “woman” (base), $c^{+}$ = “old woman” (positive), and $c^{-}$ = “young woman” (negative), enabling attribution of “old” specifically within the context of the base noun.
@@
Note that this loss is equivalent to the concept slider loss [12], which discovers concept directions in parameter space. Here, the guidance term $\left[\epsilon_{\theta}(x_{t},t;c^{+})-\epsilon_{\theta}(x_{t},t;c^{-})\right]$ corresponds to $-\nabla_{x_{t}}R(x_{t})$ in our general formulation.
@@ 3.3 Additional Techniques for Enhancing Concept-Level Attribution
To eliminate stochasticity from the forward diffusion process $x_{t}\sim q(x_{t}\mid x_{0})$, we employ deterministic DDIM inversion [41] to derive a deterministic noisy latent $x_{t}^{i}=\text{DDIMinv}(x_{0}^{i},0\rightarrow t)$ from the train sample $x_{0}^{i}$. Combined with our DPS-Loss, this approach removes all sources of randomness from gradient computation, resulting in more robust influence estimates through improved gradient fidelity. Note that this approach is limited to measuring training gradients. When measuring utility gradients, our goal is to perform data attribution for concepts in any generation process, so we create $x_{t}$ through DDIM sampling starting from $x_{T}\sim\mathcal{N}(0,I)$.
@@
As observed by Xie et al. [51], varying loss magnitudes across timesteps can cause certain gradients to dominate attribution results. To address this, we normalize each timestep gradient $g_{t}$ to unit norm, $\bar{g}_{t}=g_{t}/\|g_{t}\|_{2}$, ensuring that no single timestep exerts disproportionate influence on the final attribution score. Notably, this normalization makes our method invariant to hyperparameters such as $\beta$ and $w$ in the DPS loss, providing additional robustness.
@@ 4 Geometric Interpretation of Concept-TRAK
To understand this, consider the latent space geometry illustrated in Figure 2. In diffusion models, the noise-perturbed data at each timestep $t$ can be viewed as concentrated around a manifold (shell) $\mathcal{M}_{t}$ with high probability [8, 7]. The score function $\nabla_{x_{t}}\log p_{t}(x)$ (light orange arrow) primarily works to project noisy samples back toward the clean data manifold $\mathcal{M}_{0}$.
The key insight in our approach concerns how different loss functions interact with this manifold structure. As shown in Figure 2(a), the standard DSM loss involves the difference between unconditional and conditional score functions, i.e., $\delta_{\text{DSM}}$, which combines both projection components (pointing toward the clean data manifold) and tangential components (operating within the current manifold level). In contrast, our DPS-Loss (Figure 2(b)) isolates primarily the tangential component through the gradient $\delta_{\text{DPS}}=-\nabla_{x_{t}}\|x_{0}^{i}-\hat{x}_{0}^{i}\|^{2}$ [8, 7]. Similarly, our Reward-DPS-Loss (Figure 2(c)) creates a tangential guidance component through the reward gradient $\delta_{\text{Reward}}=\nabla_{x_{t}}R(x_{t})$.
This isolation of tangential components is crucial because they encode rich semantic information. Figure 4 demonstrates this empirically - both DPS and Reward-DPS guidance vectors maintain clean, consistent semantic structure across different timesteps of the diffusion process. This aligns with recent research showing that directions in the tangent space of diffusion models naturally correspond to interpretable concepts [30, 3, 49, 24].

### v2

@@ 1 Introduction
Data attribution methods (Koh and Liang, 2017; Park et al., 2023a; Ghorbani and Zou, 2019), which estimate how much each training example contributes to a generated output, have become valuable tools for tasks such as data valuation (Jia et al., 2019a), curation (Min et al., 2025) and understanding model behavior (Ruis et al., 2024). The impressive success of large-scale diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2021b; Ho et al., 2020) in image generation, often coupled with substantial commercial impact (Podell et al., 2023; Esser et al., 2024; Rombach et al., 2022), has been largely driven by access to massive training datasets (Schuhmann et al., 2022). However, this dependence raises pressing societal concerns, especially around copyright and data ownership (Wen et al., 2024; Somepalli et al., 2023b; Carlini et al., 2023; Somepalli et al., 2023a; GenLaw, 2024; Brittain, 2023). A key question is whether generated outputs can be traced back to specific training samples, enabling the detection of copyrighted content used during training—an area where data attribution methods show significant promise (Deng et al., 2023).
While recent work has begun exploring attribution methods tailored for diffusion models (Zheng et al., 2024; Lin et al., 2025; Mlodozeniec et al., 2024; Wang et al., 2024b), these approaches generally estimate contributions at the level of entire images. This broad perspective poses a critical limitation: in many practical scenarios, stakeholders care about specific concepts within an image, rather than the whole composition.
@@
To address this, we introduce the notion of concept-level attribution, which estimates each training example’s contribution to specific semantic features such as styles, objects, or concepts. Building on this, we propose Concept-TRAK, an extension of influence functions (Koh and Liang, 2017), to quantify how training data affects the model’s output probabilities for individual concepts. Concept-TRAK incorporates two key innovations: (1) a reformulated diffusion training loss based on diffusion posterior sampling (DPS), yielding robust, sample-specific attribution gradients; and (2) a concept-aware reward function that emphasizes semantic relevance. These components significantly enhance attribution fidelity. As shown in Figure 1(b), Concept-TRAK correctly identifies training samples responsible for the concept of Pikachu, rather than irrelevant stylistic cues—enabling precise, interpretable, and stakeholder-relevant attribution for tasks such as copyright auditing and model transparency.
To rigorously evaluate our method, we adopt a retrieval-based benchmark, AbC benchmark (Wang et al., 2023b), for concept attribution. In this setting, Concept-TRAK significantly outperforms prior methods by accurately retrieving training examples that influence specific concepts. Moreover, we demonstrate its practical value through case studies that: (1) identify training samples that contributed most to learning IP-protected concepts, (2) detect the sources of learning unsafe concepts, (3) trace the origins of both desirable features and problematic outputs for model debugging, and (4) provide insights into the origins of relational knowledge (e.g., shake hands, hugs) in diffusion models. Together, these results highlight how concept-level attribution offers valuable insights both for practical applications and for advancing our understanding of how generative models learn and combine concepts from training data.
@@ 2.1 Diffusion Model
Diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2021b; Ho et al., 2020) are a class of generative models that synthesize images through an iterative denoising process. Starting from a clean image $x_{0}$, the forward process adds Gaussian noise to produce a sequence of increasingly noisy images $x_{t}$, following, $q(x_{t}\mid x_{0})=\mathcal{N}(\sqrt{\alpha_{t}}x_{0},(1-\alpha_{t})I)$, where $\alpha_{t}$ is a noise schedule controlling the level of corruption at timestep $t$.
@@
Diffusion models can be extended to conditional generation by incorporating additional information $c$, such as a text prompt (Ho and Salimans, 2022; Nichol et al., 2021). In this setting, the model learns $\epsilon_{\theta}(x_{t},t;c)\propto\nabla\log p_{t}(x_{t}\mid c)$, allowing it to generate images that are not only realistic but also aligned with the conditioning input.
@@ 2.2 Data Attribution
The goal of data attribution is to estimate the contribution of a training sample $x^{i}$ to a model’s utility loss $\mathcal{V}$—a performance metric or objective function that quantifies how well the model performs (e.g., test loss) (Koh and Liang, 2017; Park et al., 2023a; Ghorbani and Zou, 2019). A straightforward approach is the Leave-One-Out method, which involves removing a training example $x^{i}$, retraining the model, and measuring the change in performance. However, this naive strategy is computationally expensive for modern models and large-scale datasets, as it requires retraining the model once for every training point, making it practically infeasible.
To address this limitation, influence functions (Koh and Liang, 2017) efficiently approximate the effect of removing a training example $x^{i}$ by using gradient-based estimates, avoiding the need for retraining. Given a model with parameters $\theta\in\mathbb{R}^{d}$ and training loss $\mathcal{L}(\cdot;\theta)$, the influence function is defined as:
@@
Due to the high cost of computing the full Hessian, influence function methods typically rely on inverse Hessian-vector products or approximations such as low-rank factorizations (e.g., EK-FAC) (Grosse et al., 2023). Despite these advances, influence functions remain challenging to apply in practice. Identifying training samples that strongly impact a target utility (Choe et al., 2024) requires computing gradients for every sample—an effort equivalent to a full training epoch. Unless these gradients are precomputed and stored, each attribution query results in prohibitive computational cost—especially for today’s large-scale pretrained models trained on massive datasets at great expense. Moreover, storing the full gradient for every training sample is infeasible, since each gradient matches the dimensionality of the model parameters.
To address this, TRAK (Park et al., 2023a) proposes projecting the gradients into a lower-dimensional space using a random projection matrix $P\in\mathbb{R}^{d\times k}$ with $k\ll d$, and then computing influence in the projected space:
@@ 2.3 Data attribution for diffusion models
Prior work on diffusion models (Xie et al., 2024; Zheng et al., 2024; Lin et al., 2025) has primarily focused on whole-image attribution, typically using the same objective for both training and utility losses. These studies reveal that attribution performance is highly sensitive to the choice of loss function. In particular, the standard DSM loss introduces stochasticity via both the noise term $\epsilon$ and the perturbed input $x_{t}$, resulting in noisy gradient estimates that require extensive averaging to be reliable—making it suboptimal for attribution. To mitigate this, D-TRAK (Zheng et al., 2024) employs the squared $\ell_{2}$-norm $\|\epsilon_{\theta}\|_{2}^{2}$ to compute influence scores, achieving improved stability and accuracy. However, D-TRAK does not correspond to a proper learning objective tailored to a specific sample $x^{i}_{0}$. Consequently, the resulting gradients may fail to faithfully capture the unique contribution of $x^{i}_{0}$ to the model’s learned behavior.
@@ 3.1 Train Gradients: DPS-Loss
To address the challenges associated with the stochasticity of the DSM loss and the absence of sample-specific attribution for $x^{i}_{0}$, as discussed in Section 2.3, we introduce a novel training objective grounded in the principles of diffusion posterior sampling (DPS) (Chung et al., 2022a), denoted by $\mathcal{L}_{\text{DPS}}$. This objective produces sample-specific training gradients without relying on a fully stochastic regression target and is explicitly designed to capture the influence of upweighting an individual sample $x^{i}_{0}$ on the model’s output.
@@
Consider the effect of finetuning a pretrained diffusion model $\epsilon_{\theta}$ on a single sample $x^{i}_{0}$. To model this process, we adopt a generalized Bayesian update with power likelihood where the influence of $x^{i}_{0}$ is amplified through multiplicative weighting (Holmes and Walker, 2017) (Appendix B):
@@
Instead of directly training on $x^{i}_{0}$, could we learn to model $\tilde{p}(x)$? One approach would be to use explicit score matching (ESM) (Vincent, 2011). If we had access to a diffusion model $\tilde{\epsilon}$ that already trained on $\tilde{p}(x)$, we could replace DSM with ESM for model training (Huang et al., 2021):
@@
A key advantage of ESM is its reduced reliance on the Monte Carlo-estimated regression target used in the DSM loss for influence function computation, resulting in more robust gradient estimates. Moreover, as we will show, $\tilde{\epsilon}$ can incorporate $x^{i}_{0}$ as guidance information—unlike prior approaches (Zheng et al., 2024). However, directly obtaining $\tilde{\epsilon}$ would require expensive retraining for each sample $x^{i}_{0}$, which is impractical. To address this, we introduce an efficient, feasible, and robust approximation to $\tilde{\epsilon}$.
@@
The core idea of our method is to approximate $\tilde{\epsilon}$ without additional training by using diffusion posterior sampling (DPS) principles (Chung et al., 2022a). For our target distribution $\tilde{p}(x)\propto p(x)\cdot p(x^{i}_{0}\mid x)^{w}$ , the score function at timestep $t$ becomes:
@@
[…]
@@ 3.2 Utility Gradient: Reward-DPS-Loss
Specifically, we leverage the insight that training a model with a reward function $R(x)$ modifies the learned distribution to $\tilde{p}(x)=\frac{1}{Z}p(x)\exp(\frac{1}{\beta}R(x))$, where $Z$ is a normalizing constant and $\beta$ controls the KL-divergence between $p$ and $\tilde{p}$ (Rafailov et al., 2023). For this distribution, the score function at timestep $t$ becomes:
@@
[…]
@@
We focus on measuring utility gradients for textual concepts. Extensions to other reward types, including implicit rewards (Rafailov et al., 2023), explicit reward models, and rewards for local concept attribution, are discussed in Appendix C.
@@
For textual concepts, we model the reward using conditional probabilities: $R(x_{t})\propto\log p_{t}(c\mid x_{t})$. The gradient of this reward becomes $\nabla_{x_{t}}R(x_{t})\propto\nabla_{x_{t}}\log p_{t}(x_{t}\mid c)-\nabla_{x_{t}}\log p_{t}(x_{t})$, which directly corresponds to the classifier-free guidance (Ho and Salimans, 2022).
To effectively isolate specific textual concepts, recent research (Gandikota et al., 2024; Wang et al., 2023c; Brack et al., 2023; Liu et al., 2022) demonstrates the importance of contrastive prompt engineering. Rather than using a single concept descriptor, defining concepts through comparison between positive and negative expressions yields more precise concept isolation. For example, attributing the concept “old” is more effective by comparing “old woman” to “young woman,” as this isolates age within a consistent context. Thus, we define our concept trio: $c$ = “woman” (base), $c^{+}$ = “old woman” (positive), and $c^{-}$ = “young woman” (negative), enabling attribution of “old” specifically within the context of the base noun.
@@
Note that this loss is equivalent to the concept slider loss (Gandikota et al., 2024), which discovers concept directions in parameter space. Here, the guidance term $\left[\epsilon_{\theta}(x_{t},t;c^{+})-\epsilon_{\theta}(x_{t},t;c^{-})\right]$ corresponds to $-\nabla_{x_{t}}R(x_{t})$ in our general formulation.
@@ 3.3 Additional Techniques for Enhancing Concept-Level Attribution
To eliminate stochasticity from the forward diffusion process $x_{t}\sim q(x_{t}\mid x_{0})$, we employ deterministic DDIM inversion (Song et al., 2021a) to derive a deterministic noisy latent $x_{t}^{i}=\text{DDIMinv}(x_{0}^{i},0\rightarrow t)$ from the train sample $x_{0}^{i}$. Combined with our DPS-Loss, this approach removes all sources of randomness from gradient computation, resulting in more robust influence estimates through improved gradient fidelity. Note that this approach is limited to measuring training gradients. When measuring utility gradients, our goal is to perform data attribution for concepts in any generation process, so we create $x_{t}$ through DDIM sampling starting from $x_{T}\sim\mathcal{N}(0,I)$.
@@
As observed by Xie et al. (2024), varying loss magnitudes across timesteps can cause certain gradients to dominate attribution results. To address this, we normalize each timestep gradient $g_{t}$ to unit norm, $\bar{g}_{t}=g_{t}/\|g_{t}\|_{2}$, ensuring that no single timestep exerts disproportionate influence on the final attribution score. Notably, this normalization makes our method invariant to hyperparameters such as $\beta$ and $w$ in the DPS loss, providing additional robustness.
@@ 4 Geometric Interpretation of Concept-TRAK
To understand this, consider the latent space geometry illustrated in Figure 2. In diffusion models, the noise-perturbed data at each timestep $t$ can be viewed as concentrated around a manifold (shell) $\mathcal{M}_{t}$ with high probability (Chung et al., 2022b, a). The score function $\nabla_{x_{t}}\log p_{t}(x)$ (light orange arrow) primarily works to project noisy samples back toward the clean data manifold $\mathcal{M}_{0}$.
The key insight in our approach concerns how different loss functions interact with this manifold structure. As shown in Figure 2(a), the standard DSM loss involves the difference between unconditional and conditional score functions, i.e., $\delta_{\text{DSM}}$, which combines both projection components (pointing toward the clean data manifold) and tangential components (operating within the current manifold level). In contrast, our DPS-Loss (Figure 2(b)) isolates primarily the tangential component through the gradient $\delta_{\text{DPS}}=-\nabla_{x_{t}}\|x_{0}^{i}-\hat{x}_{0}^{i}\|^{2}$ (Chung et al., 2022b, a). Similarly, our Reward-DPS-Loss (Figure 2(c)) creates a tangential guidance component through the reward gradient $\delta_{\text{Reward}}=\nabla_{x_{t}}R(x_{t})$.
This isolation of tangential components is crucial because they encode rich semantic information. Figure 4 demonstrates this empirically - both DPS and Reward-DPS guidance vectors maintain clean, consistent semantic structure across different timesteps of the diffusion process. This aligns with recent research showing that directions in the tangent space of diffusion models naturally correspond to interpretable concepts (Park et al., 2023b; Brack et al., 2023; Wang et al., 2023c; Liu et al., 2022).